# Wenjun Huang1  Jianguo Hu1

### v2

@@ 2.4 Mamba in the field of vision
The successful application of Mamba in natural language processing (NLP) has inspired its adoption in visual applications. Vision Mamba (Vim) [55] utilizes Vim blocks composed of pure Mamba layers: each Vim block models bidirectional representations using forward and backward scanning, and alleviates direction sensitivity issues in Mamba. Another approach, VMamba [34] utilizes Visual State Space (VSS) blocks that integrate Mamba and 2D convolutional layers, supported by a pyramid architecture similar to Swin Transformer [35]: each VSS block first models 2D local information through 2D deep convolution as a token mixer, and then processes 2D global information horizontally and vertically through a cross scan module. Mamba ND [27] further extends the functionality of Mamba to multidimensional data including images and videos. LocalMamba [19] segments the input image into multiple local windows and executes a state space model (SSM) in various directions within these windows to enhance local processing capabilities. EfficientVMamba [41] introduced an efficient 2D scanning technique that reduces computational requirements by performing atrous sampling on feature map blocks. In addition to these newly designed Mamba architectures, our work also draws inspiration from VL-Mamba [42], a multimodal large language model based on state space models, which has shown great potential for long-sequence modeling with fast inference and linear scaling in sequence length. Compared with these newly designed Mamba architectures, our architecture closely follows Mamba’s design ideas in the field of vision, enhancing the extraction of visual features with the latest Mamba-2 module. Our main goal with the Mamba-2 based architecture is to enhance multimodal representation and inference capabilities.
@@ 3.1 Mamba Preliminaries
The Mamba architecture inherits from state space sequence models [15], which models a $1$-D function or sequence $x(t)\in\mathbb{R}\to y(t)\in\mathbb{R}$ at time $t$ via expanded hidden states $h_{t}\in\mathbb{R}^{N}$. The hidden state is evolved through time driven by parameters $\mathbf{A},\mathbf{B},\mathbf{C}$ following linear ordinary differential equations (ODEs):
@@
By applying such transformation, we can rewrite Eq. 1 as:
@@
We then employ a matrix $\overline{\textbf{K}}$ for fast computation:
@@
where $k\in[0,L)$ and $L$ is the input sequence length. We also have $\textbf{y}=\{y_{1},...,y_{L}\}$, $\textbf{x}=\{x_{1},...,x_{L}\}$, while $\overline{\textbf{K}}\in\mathbb{R}^{L}$ can be regarded as the convolutional kernel. Note this computing structure allows Mamba to model the input sequence that perfectly matches the unidirectional, next-word prediction in autoregressive modeling.

### v3

# Wenjun Huang1,∗  Jiakai Pan1,∗  Jiahao Tang1  Yanyu Ding2 Yifei Xing3  Yuhe Wang1  Zhengzhuo Wang1  Jianguo Hu1,†

@@ 2.4 Mamba in the field of vision
The successful application of Mamba in natural language processing (NLP) has inspired its adoption in visual applications [43]. Vision Mamba (Vim) [56] utilizes Vim blocks composed of pure Mamba layers: each Vim block models bidirectional representations using forward and backward scanning, and alleviates direction sensitivity issues in Mamba. Another approach, VMamba [34] utilizes Visual State Space (VSS) blocks that integrate Mamba and 2D convolutional layers, supported by a pyramid architecture similar to Swin Transformer [35]: each VSS block first models 2D local information through 2D deep convolution as a token mixer, and then processes 2D global information horizontally and vertically through a cross scan module. Mamba ND [27] further extends the functionality of Mamba to multidimensional data including images and videos. LocalMamba [19] segments the input image into multiple local windows and executes a state space model (SSM) in various directions within these windows to enhance local processing capabilities. EfficientVMamba [41] introduced an efficient 2D scanning technique that reduces computational requirements by performing atrous sampling on feature map blocks. In addition to these newly designed Mamba architectures, our work also draws inspiration from VL-Mamba [42], a multimodal large language model based on state space models, which has shown great potential for long-sequence modeling with fast inference and linear scaling in sequence length. Compared with these newly designed Mamba architectures, our architecture closely follows Mamba’s design ideas in the field of vision, enhancing the extraction of visual features with the latest Mamba-2 module. Our main goal with the Mamba-2 based architecture is to enhance multimodal representation and inference capabilities.
@@ 3.1 Mamba Preliminaries
The Mamba architecture is derived from state space sequence models [15], which models a $1$-D function or sequence $x(t)\in\mathbb{R}\to y(t)\in\mathbb{R}$ at time $t$ via expanded hidden states $h_{t}\in\mathbb{R}^{N}$. These hidden states evolve over time according to parameters $\mathbf{A},\mathbf{B},\mathbf{C}$ and are governed by linear ordinary differential equations (ODEs):
@@
Using this transformation, Eq. 1 can be rewritten as:
@@
We then utilize the matrix $\overline{\textbf{K}}$ to enable efficient computation:
@@
where $k\in[0,L)$ and $L$ is the input sequence length. We also have $\textbf{y}=\{y_{1},...,y_{L}\}$, $\textbf{x}=\{x_{1},...,x_{L}\}$, while $\overline{\textbf{K}}\in\mathbb{R}^{L}$ can be regarded as the convolutional kernel.